# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 3 The greatest possible dimension for a trivial spectrum subspace
This short section is devoted to the proof of Theorem 1.6. We have seen right at the beginning of the article that for every integer $n\geq 1$ there exists an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $\alpha(n,d)$. Hence, in order to prove Theorem 1.6 it suffices to prove the following result:
@@
Let $\mathbb{D}$ be a division ring with finite degree $d$ over a central subfield $\mathbb{F}$, and $V$ be a vector space over $\mathbb{D}$ with finite dimension $n$. Assume that $|\mathbb{F}|\geq dn$. Let $\mathcal{S}$ be an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Then
$\dim\mathcal{S}\leq\alpha(n,d)=n(d-1)+d\dbinom{n}{2}$.
The main key to this result is the use of the dual operator space of $\mathcal{S}$ (see Section 1.5) and what we call the Tangent Space Lemma. The latter is recalled in the next section.
@@ 3.1 The tangent space lemma, and the Flanders-Atkinson lemma
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{F}$, and $\mathcal{S}$ be a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$. In $\mathcal{S}$, take an element $u_{0}$ of maximal rank $r$, and assume that $|\mathbb{F}|>r$. Then every element of $\mathcal{S}$ maps $\operatorname{\operatorname{Ker}}u_{0}$ into $\operatorname{\operatorname{Im}}u_{0}$.
@@
We prove the result by induction on $n$. The case $n=0$ is trivial, so we assume that $n\geq 1$. Let us consider the dual operator space $\widehat{\mathcal{S}}$, consisting of the $\mathbb{F}$-linear mappings
$\widehat{x}:u\in\mathcal{S}\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
Note that $\widehat{\mathcal{S}}$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S},V)$. Let $x\in V\smallsetminus\{0\}$. Since $\mathcal{S}$ has $\mathbb{F}$-trivial spectrum, there is no $u\in\mathcal{S}$ such that $u(x)=x$; hence $\operatorname{\operatorname{rk}}(\widehat{x})\leq dn-1$.
Let us now take $x\in V\smallsetminus\{0\}$ such that $\widehat{x}$ has the greatest possible rank $r\leq dn-1$ in $\widehat{\mathcal{S}}$. Since $|\mathbb{F}|>r$, the Tangent Space Lemma (Lemma 3.3) yields
@@
$\forall u\in\mathcal{S},\;u(x)=0\Rightarrow\operatorname{\operatorname{Im}}u\subseteq\mathcal{S}x.$
@@
$\mathcal{S}^{\prime}:=\operatorname{\operatorname{Ker}}\widehat{x}=\{u\in\mathcal{S}:\;u(x)=0\},$
@@
$\dim(\mathcal{S})=\dim(\mathcal{S}^{\prime})+\dim_{\mathbb{F}}(\mathcal{S}x)\leq\dim(\mathcal{S}^{\prime})+(dn-1).$ (4)
Now, the operators in $\mathcal{S}^{\prime}$ are actually $\mathbb{D}$-linear so their ranges are $\mathbb{D}$-linear subspaces of $V$, all included in $\mathcal{S}x$ and hence all included in the socle (see Definition 2.1) of $\mathcal{S}x$. This socle does not contain $x$ (because $x\not\in\mathcal{S}x$), and hence we can embed it into a $\mathbb{D}$-hyperplane $H$ of $V$ such that $x\mathbb{D}\oplus H=V$, and we note that $\operatorname{\operatorname{Im}}u\subseteq H$ for every $u\in\mathcal{S}^{\prime}$.
Hence, all the elements of $\mathcal{S}^{\prime}$ leave $H$ invariant, inducing $\mathbb{D}$-linear endomorphisms of $H$, all of them with $\mathbb{F}$-trivial spectrum. Since $x\mathbb{D}\oplus H=V$ and the elements of $\mathcal{S}^{\prime}$ vanish on $x\mathbb{D}$, the $\mathbb{F}$-linear mapping
$u\in\mathcal{S}^{\prime}\mapsto u_{H}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$
is injective. Its range is an $\mathbb{F}$-trivial spectrum subspace $\mathcal{T}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$. Since $|\mathbb{F}|\geq(n-1)d$, we obtain by induction that
$\dim(\mathcal{S}^{\prime})=\dim(\mathcal{T})\leq d\dbinom{n-1}{2}+(n-1)(d-1).$
@@ 3.3 Adaptation to deeply intransitive operator spaces
Let $U$ and $V$ be $\mathbb{D}$-vector spaces, with $n=\dim_{\mathbb{D}}V\geq 1$. Assume that $|\mathbb{F}|\geq nd$. Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Then $\dim\mathcal{S}\leq\alpha(n,d)$.
@@
We denote by $r$ the transitive rank of $\mathcal{S}$. If $r=0$ then $\mathcal{S}=\{0\}$ and we are done. Assume that $r>0$, and choose $x\in U$ such that $\dim_{\mathbb{F}}\mathcal{S}x=r$. Hence $\widehat{x}$ has the greatest possible rank in $\widehat{\mathcal{S}}$, and as $|\mathbb{F}|>r$ we can apply the tangent space lemma. This yields that all the operators in $\mathcal{S}^{\prime}:=\{f\in\mathcal{S}:f(x)=0\}$ map into $\mathcal{S}x$. As in the previous proof, we can find a $\mathbb{D}$-linear hyperplane $H$ of $V$ into which all the elements of $\mathcal{S}^{\prime}$ map. The rank theorem yields
$\dim\mathcal{S}=r+\dim\mathcal{S}^{\prime}.$
@@
  • If $n=1$ then $\mathcal{S}^{\prime}=\{0\}$ and we conclude that $\dim\mathcal{S}\leq d-1$, as claimed;
  • If $n>1$ then $\mathcal{S}^{\prime}$ is a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, by induction we find that
    $\dim\mathcal{S}\leq(nd-1)+\alpha(n-1,d)=\alpha(n,d),$
@@
Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(3-d,1)$ and $n>0$. Then $\mathcal{S}$ is target-reduced.
@@
In the proof and in future use of the result, it is crucial to note that the fact that $\mathcal{S}$ is not target-reduced is equivalent to the existence of a $\mathbb{D}$-linear hyperplane $H$ of $V$ (and not simply of an $\mathbb{F}$-linear hyperplane) that includes the range of every element of $\mathcal{S}$. The converse implication is obvious. For the direct one, assume that we have an $\mathbb{F}$-linear hyperplane $V^{\prime}$ of $V$ that includes the range of every operator in $\mathcal{S}$. Then all the operators in $\mathcal{S}$ map into the socle of $V^{\prime}$, which is a $\mathbb{D}$-linear hyperplane of $V$.
@@
Assume that $\mathcal{S}$ is not target-reduced. Then there exists a $\mathbb{D}$-linear hyperplane $H$ of $V$ such that $\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
Because $\mathcal{S}$ is deeply intransitive we can view it as a deeply intransitive subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, and then Proposition 3.4 yields $\dim\mathcal{S}\leq\alpha(n-1,d)=\alpha(n,d)-(nd-1)$. Now (nd1)n=n(d1)1(nd-1)-n=n(d-1)-1, and n(d1)1>2n(d-1)-1>-2 if $d=1$, whereas n(d1)1>1n(d-1)-1>-1 whenever $d\geq 2$ and $n>0$. Hence, in any case we contradict the asumption that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(3-d,1)$. ∎
@@ 4.3 Recognizing alternators in generic matrix identities
Let us take finite-dimensional vector spaces $U$ and $V$ over $\mathbb{F}$, a linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, and let $\mathbf{B}_{\mathcal{S}}=(v_{1},\dots,v_{n})$ and $\mathbf{B}_{V}=(e_{1},\dots,e_{p})$ be respective bases of $\mathcal{S}$ and $V$.
Throughout, we assume that $\mathcal{S}$ is target-reduced (see the definition in Section 1.5), to the effect that $\widehat{\mathcal{S}}$ is also target-reduced. We consider the matrix space $\mathcal{M}$ that represents the dual operator space $\widehat{\mathcal{S}}$ in $\mathbf{B}_{\mathcal{S}}$ and $\mathbf{B}_{V}$, and we take an arbitrary generic matrix $\mathbf{M}$ that is associated with $\mathcal{M}$, with parameter space denoted by $E$. We define the following linear mappings:
  • $\Phi$ assigns to every $z\in E$ the operator in $\widehat{\mathcal{S}}$ whose matrix in the bases $\mathbf{B}_{\mathcal{S}}$ and $\mathbf{B}_{V}$ is $\mathbf{M}(z)$;
  • $\widehat{(-)}$ assigns to every $y\in U$ the evaluation mapping $\widehat{y}\in\widehat{\mathcal{S}}$.
@@
Now, we consider an alternator $b$ of $\mathcal{S}$. The left mapping $L_{b}:y\in U\mapsto b(y,-)\in V^{\star}$ can then be inserted into the following diagram:
@@
Next, we prove that $L_{b}$ factors into the composite of $\widehat{(-)}$ with a unique linear mapping $\Theta:\widehat{\mathcal{S}}\rightarrow V^{\star}$. To see this, we apply the factorization lemma for linear maps: all we need is to check that $L_{b}$ vanishes at every vector $y\in U$ such that $\widehat{y}=0$. So, let $y\in U$ satisfy $\widehat{y}=0$. Hence, for all $f\in\mathcal{S}$, we have b(y,f(z))=b(z,f(y))=0b(y,f(z))=-b(z,f(y))=0 because $f$ is $b$-alternating. Because $\mathcal{S}$ is target-reduced, this yields $y\in\operatorname{\operatorname{Ker}}L_{b}$, as claimed.
@@
$\Lambda:\operatorname{\operatorname{Alt}}(\mathcal{S})\hookrightarrow\operatorname{\operatorname{catch}}(\mathbf{M}).$
Conversely, let $\mathbf{L}$ be a catcher of $\mathbf{M}$. Let $\Psi:E\rightarrow V^{\star}$ assign to every $z\in E$ the linear form on $V$ whose matrix in the basis $\mathbf{B}_{V}$ is $\mathbf{L}(z)$. We claim that there is a (unique) factorization $\Psi=\Theta\circ\Phi$ for some linear mapping $\Theta:\widehat{\mathcal{S}}\rightarrow V^{\star}$. Again, we use the factorization lemma. Let $z_{0}\in\operatorname{\operatorname{Ker}}\Phi$. Let us polarize the quadratic identity $\mathbf{L}\mathbf{M}=0$ at $z_{0}$. Because $\mathbf{M}(z_{0})=0$ by assumption, this yields $\mathbf{L}(z_{0})M=0$ for all $M\in\mathcal{M}$. In other words, every operator in $\widehat{\mathcal{S}}$ maps into $\operatorname{\operatorname{Ker}}\Psi(z_{0})$. Since $\widehat{\mathcal{S}}$ is target-reduced, we deduce that $\Psi(z_{0})=0$. Hence the existence of $\Theta$, the uniqueness of which is obvious because $\Phi$ is onto. Finally, we consider the composite mapping

### v2

@@ 3 The greatest possible dimension for a trivial spectrum subspace
This short section is devoted to the proof of Theorem 1.8. We have seen right at the beginning of Section 1.2 that for every integer $n\geq 1$ there exists a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $\alpha(n,d)=n(d-1)+d\dbinom{n}{2}$. Hence, in order to prove Theorem 1.8 it suffices to prove the following result:
@@
Let $\mathbb{D}$ be a division ring with finite dimension $d$ over a central subfield $\mathbb{F}$, and $V$ be a vector space over $\mathbb{D}$ with finite dimension $n$. Assume that $|\mathbb{F}|\geq dn$. Let $S$ be a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Then
$\dim S\leq\alpha(n,d)=n(d-1)+d\dbinom{n}{2}$.
The main key to this result is the use of the dual operator space of $S$ (see Section 1.5) and what we call the Tangent Space Lemma. The latter is recalled in the next section.
@@ 3.1 The Tangent Space Lemma, and the Flanders-Atkinson lemma
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{F}$, and $S$ be a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$. In $S$, take an element $u_{0}$ of maximal rank $r$, and assume that $|\mathbb{F}|>r$. Then every element of $S$ maps $\operatorname{\operatorname{Ker}}u_{0}$ into $\operatorname{\operatorname{Im}}u_{0}$.
@@
We prove the result by induction on $n$. The case $n=0$ is trivial, so we assume that $n\geq 1$. Let us consider the dual operator space $\widehat{S}$, consisting of the $\mathbb{F}$-linear mappings
$\widehat{x}:u\in S\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
Note that $\widehat{S}$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(S,V)$. Let $x\in V\smallsetminus\{0\}$. Since $S$ has trivial spectrum, there is no $u\in S$ such that $u(x)=x$; hence $\operatorname{\operatorname{rk}}(\widehat{x})\leq dn-1$.
Let us now take $x\in V\smallsetminus\{0\}$ such that $\widehat{x}$ has the greatest possible rank $r\leq dn-1$ in $\widehat{S}$. Since $|\mathbb{F}|>r$, the Tangent Space Lemma (Lemma 3.3) yields
@@
$\forall u\in S,\;u(x)=0\Rightarrow\operatorname{\operatorname{Im}}u\subseteq Sx.$
@@
$S^{\prime}:=\operatorname{\operatorname{Ker}}\widehat{x}=\{u\in S:\;u(x)=0\},$
@@
$\dim(S)=\dim(S^{\prime})+\dim_{\mathbb{F}}(Sx)\leq\dim(S^{\prime})+(dn-1).$ (2)
Now, the operators in $S^{\prime}$ are actually $\mathbb{D}$-linear so their ranges are $\mathbb{D}$-linear subspaces of $V$, all included in $Sx$ and hence all included in the socle (see Definition 2.1) of $Sx$. This socle does not contain $x$ (because $x\not\in Sx$), and hence we can embed it into a $\mathbb{D}$-hyperplane $H$ of $V$ such that $x\mathbb{D}\oplus H=V$, and we note that $\operatorname{\operatorname{Im}}u\subseteq H$ for every $u\in S^{\prime}$.
Hence, all the elements of $S^{\prime}$ leave $H$ invariant, inducing $\mathbb{D}$-linear endomorphisms of $H$, all of them with trivial spectrum. Since $x\mathbb{D}\oplus H=V$ and the elements of $S^{\prime}$ vanish on $x\mathbb{D}$, the $\mathbb{F}$-linear mapping
$u\in S^{\prime}\mapsto u_{H}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$
is injective. Its range is a trivial spectrum subspace $\mathcal{T}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$. Since $|\mathbb{F}|\geq(n-1)d$, we obtain by induction that
$\dim(S^{\prime})=\dim(\mathcal{T})\leq d\dbinom{n-1}{2}+(n-1)(d-1).$
@@ 3.3 Adaptation to deeply intransitive operator spaces
Let $U$ and $V$ be $\mathbb{D}$-vector spaces, with $n=\dim_{\mathbb{D}}V\geq 1$. Assume that $|\mathbb{F}|\geq nd$. Let $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Then $\dim S\leq\alpha(n,d)$.
@@
We denote by $r$ the transitive rank of $S$. If $r=0$ then $S=\{0\}$ and we are done. Assume that $r>0$, and choose $x\in U$ such that $\dim_{\mathbb{F}}Sx=r$. Hence $\widehat{x}$ has the greatest possible rank in $\widehat{S}$, and as $|\mathbb{F}|>r$ we can apply the tangent space lemma. This yields that all the operators in $S^{\prime}:=\{f\in S:f(x)=0\}$ map into $Sx$. As in the previous proof, we can find a $\mathbb{D}$-linear hyperplane $H$ of $V$ into which all the elements of $S^{\prime}$ map. The rank theorem yields
$\dim S=r+\dim S^{\prime}.$
@@
  • If $n=1$ then $S^{\prime}=\{0\}$ and we conclude that $\dim S\leq d-1$, as claimed;
  • If $n>1$ then $S^{\prime}$ is a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, by induction we find that
    $\dim S\leq(nd-1)+\alpha(n-1,d)=\alpha(n,d),$
@@
Let $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim S\geq\alpha(n,d)-n+\max(3-d,1)$ and $n>0$. Then $S$ is target-reduced.
@@
In the proof and in future use of the result, it is crucial to note that the fact that $S$ is not target-reduced is equivalent to the existence of a $\mathbb{D}$-linear hyperplane $H$ of $V$ (and not simply of an $\mathbb{F}$-linear hyperplane) that includes the range of every element of $S$. The converse implication is obvious. For the direct one, assume that we have an $\mathbb{F}$-linear hyperplane $V^{\prime}$ of $V$ that includes the range of every operator in $S$. Then all the operators in $S$ map into the socle of $V^{\prime}$, which is a $\mathbb{D}$-linear hyperplane of $V$.
@@
Assume that $S$ is not target-reduced. Then there exists a $\mathbb{D}$-linear hyperplane $H$ of $V$ such that $S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
Because $S$ is deeply intransitive we can view it as a deeply intransitive subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, and then Proposition 3.4 yields $\dim S\leq\alpha(n-1,d)=\alpha(n,d)-(nd-1)$. Now (nd1)n=n(d1)1(nd-1)-n=n(d-1)-1, and n(d1)1>2n(d-1)-1>-2 if $d=1$, whereas n(d1)1>1n(d-1)-1>-1 whenever $d\geq 2$ and $n>0$. Hence, in any case we contradict the asumption that $\dim S\geq\alpha(n,d)-n+\max(3-d,1)$. ∎
@@ 4.3 Recognizing alternators in generic matrix identities
Let us take finite-dimensional vector spaces $U$ and $V$ over $\mathbb{F}$, a linear subspace $S$ of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, and let $\mathbf{B}_{S}=(v_{1},\dots,v_{n})$ and $\mathbf{B}_{V}=(e_{1},\dots,e_{p})$ be respective bases of $S$ and $V$.
Throughout, we assume that $S$ is target-reduced (see the definition in Section 1.5), to the effect that $\widehat{S}$ is also target-reduced. We consider the matrix space $\mathcal{M}$ that represents the dual operator space $\widehat{S}$ in $\mathbf{B}_{S}$ and $\mathbf{B}_{V}$, and we take an arbitrary generic matrix $\mathbf{M}$ that is associated with $\mathcal{M}$, with parameter space denoted by $E$. We define the following linear mappings:
  • $\Phi$ assigns to every $z\in E$ the operator in $\widehat{S}$ whose matrix in the bases $\mathbf{B}_{S}$ and $\mathbf{B}_{V}$ is $\mathbf{M}(z)$;
  • $\widehat{(-)}$ assigns to every $y\in U$ the evaluation mapping $\widehat{y}\in\widehat{S}$.
@@
Now, we consider an alternator $b$ of $S$. The left mapping $L_{b}:y\in U\mapsto b(y,-)\in V^{\star}$ can then be inserted into the following diagram:
@@
Next, we prove that $L_{b}$ factors into the composite of $\widehat{(-)}$ with a unique linear mapping $\Theta:\widehat{S}\rightarrow V^{\star}$. To see this, we apply the factorization lemma for linear maps: all we need is to check that $L_{b}$ vanishes at every vector $y\in U$ such that $\widehat{y}=0$. So, let $y\in U$ satisfy $\widehat{y}=0$. Hence, for all $f\in S$, we have b(y,f(z))=b(z,f(y))=0b(y,f(z))=-b(z,f(y))=0 because $f$ is $b$-alternating. Because $S$ is target-reduced, this yields $y\in\operatorname{\operatorname{Ker}}L_{b}$, as claimed.
@@
$\Lambda:\operatorname{\operatorname{Alt}}(S)\hookrightarrow\operatorname{\operatorname{catch}}(\mathbf{M}).$
Conversely, let $\mathbf{L}$ be a catcher of $\mathbf{M}$. Let $\Psi:E\rightarrow V^{\star}$ assign to every $z\in E$ the linear form on $V$ whose matrix in the basis $\mathbf{B}_{V}$ is $\mathbf{L}(z)$. We claim that there is a (unique) factorization $\Psi=\Theta\circ\Phi$ for some linear mapping $\Theta:\widehat{S}\rightarrow V^{\star}$. Again, we use the factorization lemma. Let $z_{0}\in\operatorname{\operatorname{Ker}}\Phi$. Let us polarize the quadratic identity $\mathbf{L}\mathbf{M}=0$ at $z_{0}$. Because $\mathbf{M}(z_{0})=0$ by assumption, this yields $\mathbf{L}(z_{0})M=0$ for all $M\in\mathcal{M}$. In other words, every operator in $\widehat{S}$ maps into $\operatorname{\operatorname{Ker}}\Psi(z_{0})$. Since $\widehat{S}$ is target-reduced, we deduce that $\Psi(z_{0})=0$. Hence the existence of $\Theta$, the uniqueness of which is obvious because $\Phi$ is onto. Finally, we consider the composite mapping